c) Fanta ltd has a dividend cover of five (5) times and recorded the following earnings after tax; 
YEAR EARNINGS $
2011 100,000
2012 120,000
2013 180,000
2014 220,000
2015 300,000
Calculate the average dividend growth rate for Fanta ltd.
1
Expert's answer
2020-11-22T18:43:18-0500

Divided growth in 2012

"=[\frac{120000}{100000}-1]\times100\%
= 20\%"

Divided growth in 2013

"=[\frac{180000}{120000}-1]\times100\%
= 50\%"

Divided growth in 2014

"=[\frac{220000}{180000}-1]\times100\%
= 22.2\%"

Divided growth in 2015

"=[\frac{300000}{220000}-1]\times100\%
= 36.4\%"

Therefore, Average growth Rate

"n=2015-2011=4"

"=\frac{20+50+22.2+36.4}{4}=32.15\%"